# Implicit Dynamical Flow Fusion (IDFF) for Generative Modeling

### v1

@@ Abstract
Conditional Flow Matching (CFM) models can generate high-quality samples from a non-informative prior but they can be slow, often needing hundreds of network evaluations (NFE). To address this, we propose Implicit Dynamical Flow Fusion (IDFF); IDFF learns a new vector field with an additional momentum term that enables taking longer steps during sample generation while maintaining the fidelity of the generated distribution. Consequently, IDFFs reduce the NFEs by a factor of ten (relative to CFMs) without sacrificing sample quality, enabling rapid sampling and efficient handling of image and time-series data generation tasks. We evaluate IDFF on standard benchmarks such as CIFAR-10 and CelebA for image generation. We achieved comparable likelihood and quality performance to CFMs and diffusion-based models with many fewer NFEs. IDFF also shows superior performance on time-series datasets modeling, including molecular simulation and sea surface temperature (SST) datasets, highlighting its versatility and effectiveness across different domains.
@@ 1 Introduction
Diffusion models (Ho et al., 2020; Cachay et al., 2023; Myers et al., 2022) have gained significant traction in modeling complex, high-dimensional data by iteratively transforming random noise into meaningful data. These models have been successfully applied in diverse fields, such as generating high-quality images (Song et al., 2020a) and text (Liu et al., 2024; Kim et al., 2019). However, the training process is computationally intensive, requiring numerous steps of sampling (Ho et al., 2020; Song et al., 2020b). Specifically, diffusion models require hundreds of function evaluations (NFEs) to achieve high-quality results. Techniques such as DPM-solver (Lu et al., 2022a, b) and Denoising Diffusion Implicit Models (DDIMs) (Song et al., 2020b) address help reduce the NFE for diffusion-based generative models. While diffusion models parameterize the dynamics that transform noise into data, Conditional Flow Matching (CFMs) (Li et al., 2022; Tong et al., 2023; Liu et al., 2022; Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023) directly parameterize vector fields that govern transformation from noise to data. CFMs define conditional probability paths through the change of variable theory in statistics (Lipman et al., 2022), enabling faster convergence during training compared to diffusion models (Lipman et al., 2022).
Figure 1: Image generation using IDFF (with NFE=10) has been applied to multiple datasets, including CIFAR10, CelebA-64, ImageNet-64, LSUN-Bedroom, LSUN-Church, and CelebA-24. The quality of the generated images demonstrates IDFF’s ability to capture and reproduce realistic visuals across varying levels of complexity and resolution. Additional samples and further analysis can be found in Appendix A.2.
Generating high-quality samples with CFMs requires hundreds of NFEs; this makes the process computationally expensive (Dao et al., 2023; Lipman et al., 2022) and restricts the extension of CFM models to time-series data modeling since the NFE scale with the length of the sequence.
To overcome these challenges, we introduce a novel method to improve the efficiency of CFMs, allowing for longer sampling steps without compromising the fidelity of the target distribution. Our approach is inspired by Hamiltonian Monte Carlo (HMC) algorithms (Neal et al., 2011; Neal, 2012), which utilize the conservation properties of Hamiltonian dynamics to couple the target and momentum variables.
The contributions of our work are as follows:
@@
    We propose the Implicit Dynamical Flow Fusion (IDFF), which integrates a momentum term into the vector field of conditional flow models. Similar to how HMC reduces the number of samples required to reach the target distribution (Neal, 2012; Hoffman et al., 2014), IDFF reduces the NFE by a factor of 10 compared to traditional CFMs while maintaining identical marginal distributions. Additionally, we introduce a training objective that directly aligns the generated samples with the actual data rather than solely focusing on matching the conditional flows.
@@
    We validate IDFF through extensive experiments on image generation using datasets such as CIFAR-10 and CelebA. Additional results on ImageNet-64, CelebA-HQ, LSUN-Church, and LSUN-Bedroom are presented in Figure 1. Our models achieve a superior balance between computational cost and sample quality compared to both traditional diffusion models (e.g., DDPM (Ho et al., 2020), DDIM (Song et al., 2020b)) and current CFMs (Pooladian et al., 2023; Tong et al., 2023).
@@
    We show that reducing NFEs makes IDFF practical for other generation tasks, such as time-series modeling. We demonstrate IDFF’s effectiveness in capturing complex dynamical systems, including 3D attractors, molecular dynamics (MD), and sea surface temperatures (SST). Our results indicate that IDFF surpasses other methods regarding sample quality for MD and SST datasets while maintaining computational efficiency, with NFE $\leq$ 5 for each generated sample.
@@ 2 Background
Score-based generative models
@@
Score-based models (Song et al., 2020b; Song and Ermon, 2019) represent a broad class of diffusion models that describe a continuous-time stochastic process. These models are characterized by a stochastic process ${\mathbf{x}}_{t}$, which is governed by the following Ito equation:
$d{\mathbf{x}}=f({\mathbf{x}},t)dt+g(t)d{\mathbf{w}},$ (1)
@@
Where $t\in[0,1]$, $f(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, and $g(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ represent the drift and diffusion coefficients, respectively, and ${\mathbf{w}}\in\mathbb{R}^{d}$ is a standard Wiener process.
@@
The time-reversed version of this diffusion process, obtained from the Fokker–Planck equations(Song et al., 2020b), is also a diffusion process. This reverse-time stochastic SDE is defined as:
$d{\mathbf{x}}=\left(f({\mathbf{x}},t)-g(t)^{2}\nabla_{{\mathbf{x}}}\log p_{t}({\mathbf{x}})\right)dt+g(t)d\bar{{\mathbf{w}}},$ (2)
@@
where $\bar{{\mathbf{w}}}$ is a standard Wiener process. Equation 2 produces the same marginal distributions as the forward diffusion process defined by Equation 1 (Song et al., 2020b).
To sample from this SDE, one must solve Equation 2, beginning with random noise as the initial condition, as demonstrated by (Ho et al., 2020) ${\mathbf{x}}_{0}\sim\mathcal{N}(0,\mathbf{I})$, and gradually converges to the data distribution $p_{\mathrm{data}}$ by time $t=1$ (Song et al., 2020a). The score-matching approach enables this by choosing $f({\mathbf{x}},t)=h(t){\mathbf{x}}$, a fixed affine transformation of ${\mathbf{x}}$, while learning $\nabla_{{\mathbf{x}}}\log p_{t}({\mathbf{x}})$ through a score-matching loss (Ho et al., 2020).
Conditional flow matching as a generative model
A CFM model (Lipman et al., 2022) defines a time-dependent vector field ${\mathbf{v}}_{t}$ parameterized by a neural network that creates a conditional flow, $\phi_{t}(\cdot|{\mathbf{x}}_{1}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, defined by an ODE:
$\frac{d}{dt}\phi_{t}({\mathbf{x}}|{\mathbf{x}}_{1})={\mathbf{v}}_{t}(\phi_{t}({\mathbf{x}}|{\mathbf{x}}_{1}))$ (3)
with an initial condition $\phi_{0}({\mathbf{x}}_{0}|{\mathbf{x}}_{1})={\mathbf{x}}_{0}$. A CFM uses a continuous normalizing flow (CNF) to transform a simple prior density $p_{0}$ (e.g., pure noise) into a more complex one, $p_{1}$, using the push-forward equation:
$p_{t}=p_{0}(\phi_{t}^{-1}({\mathbf{x}}|{\mathbf{x}}_{1}))\left|\textbf{det}\left[\frac{\partial\phi_{t}^{-1}}{\partial{\mathbf{x}}}({\mathbf{x}}|{\mathbf{x}}_{1})\right]\right|$ (4)
Therefore, a generative conditional vector field represents a conditional probability density path $p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$, a time-dependent probability density function for $t\in[0,1]$, generated by a vector field $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$. Here, $p_{0}$ is a simple prior distribution, and $p_{1}\approx p_{data}$. Given a target conditional probability density path $p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$ and a corresponding conditional vector field $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$, which generates $p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$, we can define the CFM objective:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,{\mathbf{x}}_{1},{\mathbf{x}}_{t}}\big{\|}{\mathbf{v}}_{t}({\mathbf{x}}_{t})-\hat{{\mathbf{v}}}_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})\big{\|}^{2}$ (5)
where $t\sim{\mathcal{U}}[0,1]$, ${\mathbf{x}}_{1}\sim p_{data}$, and ${\mathbf{x}}_{t}\sim p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})$. The training objective defined in 5 guarantees that the marginalized $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$, $\hat{{\mathbf{v}}}_{t}({\mathbf{x}})$, generates $p_{t}({\mathbf{x}})$ as FM models do, and thus allows easy sampling of unbiased estimates as long as we can efficiently sample from $p_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$ and compute $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}|{\mathbf{x}}_{1})$, both of which can be done as they are defined on a per-sample basis (Lipman et al., 2022).
@@
CFMs share conceptual similarities with diffusion models, as both frameworks rely on defining continuous probability paths over time. While diffusion models use stochastic processes to transform data distributions, typically governed by an SDE as seen in (Song et al., 2020b; Ho et al., 2020), CFMs utilize a deterministic vector field to create a conditional flow that directly maps an initial prior distribution to the target distribution via an ODE. Diffusion models, such as denoising score matching models, rely on stochastic diffusion paths to approximate the data distribution. At the same time, CFMs bypass this stochastic process by constructing a conditional vector field and sampling from an ODE to generate samples (Lipman et al., 2022; Song et al., 2020a).
@@
The CFM objective defined in Equation 5 appears simple. However, it is nearly intractable due to the absence of prior knowledge regarding an appropriate coupling term between the initial ($p_{0}$) and target ($p_{1}$) distributions. To address this challenge, optimal transport (OT) (Pooladian et al., 2023; Lipman et al., 2022) is used to couple the initial and target distributions. This term seeks a mapping from $p_{0}$ to $p_{1}$ that minimizes the displacement cost between these two distributions, resulting in flows that can be integrated accurately. This CFM variant is Optimal Transport CFMs (OT-CFMs)(Tong et al., 2023). Although OT improves the efficiency of training in CFMs, as shown in Figure 2, traditional sampling methods for OT-CFMs still require more than a hundred function evaluations (NFE) to generate high-quality samples (Dao et al., 2023; Tong et al., 2023).
@@
Figure 2: A) Sampling Trajectories Comparison Between IDFF and OT-CFMs: The figure shows 4096 final samples generated by IDFF. As observed, IDFF (represented by the blue arrows) takes larger steps toward the target distribution, guided by the momentum term (shown by the red arrows). Although OT-CFM follows an almost straight path to reach the target distribution, it requires many function evaluations (NFE), whereas IDFF achieves this 20 times faster. B) IDFF Sampling Process: In this process, $\hat{{\mathbf{x}}}1(.,t;\theta)$ denotes the denoising model, and $\epsilon({\mathbf{x}},t;\theta)$ represents the scoring model. The term $\sigma_{t}=\sigma_{0}\sqrt{t(1-t)}$ controls the interaction between the flow and score models, creating the vector field ${{\mathbf{w}}_{t}}$ that generates data samples $x{t+\Delta t}\sim\mathcal{N}({\mathbf{x}}_{t}+{\mathbf{w}}_{t}\Delta t,\sigma_{t}^{2}\Delta t\textbf{I})$. For further details, see Algorithm 2.
As shown in Pooladian et al. (2023), the vector field associated with the conditional vector field $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})=\frac{{\mathbf{x}}_{1}-{\mathbf{x}}}{1-t}$ can be expressed as $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})={\mathbf{x}}_{1}-{\mathbf{x}}_{0}$. It becomes a deterministic function when conditioned on ${\mathbf{x}}_{0}$. Consequently, using any deterministic solver and conditioning on the initial points ${\mathbf{x}}_{0}$, the sample generation process becomes deterministic, which limits the generalizability of the conditional OT-CFM models compared to the stochastic processes employed in denoising diffusion models. Moreover, as flow matching learns $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})$, there is no flexibility in the sampling steps during the training phase. This restriction on sampling, constrained to the integration steps, hinders the ability of OT-CFMs to generate high-quality samples with a small number of integration steps, often leading to a high NFE during the sampling process.

### v2

@@ Abstract
Conditional Flow Matching (CFM) models can generate high-quality samples from a non-informative prior, but they can be slow, often needing hundreds of network evaluations (NFE). To address this, we propose Implicit Dynamical Flow Fusion (IDFF); IDFF learns a new vector field with an additional momentum term that enables taking longer steps during sample generation while maintaining the fidelity of the generated distribution. Consequently, IDFFs reduce the NFEs by a factor of ten (relative to CFMs) without sacrificing sample quality, enabling rapid sampling and efficient handling of image and time-series data generation tasks. We evaluate IDFF on standard benchmarks such as CIFAR-10 and CelebA for image generation, where we achieve likelihood and quality performance comparable to CFMs and diffusion-based models with fewer NFEs. IDFF also shows superior performance on time-series datasets modeling, including molecular simulation and sea surface temperature (SST) datasets, highlighting its versatility and effectiveness across different domains.
@@ 1 Introduction
Diffusion models have emerged as powerful tools for modeling complex, high-dimensional data by iteratively transforming random noise into meaningful information (Cachay et al., 2023; Myers et al., 2022). These models have demonstrated remarkable success in various domains, including high-quality image generation (Song et al., 2020a) and text production (Liu et al., 2024; Kim et al., 2019). However, their training process remains computationally intensive, often requiring hundreds of function evaluations (NFEs) to achieve high-quality results (Ho et al., 2020). While techniques such as DPM-solver (Lu et al., 2022a, b) and Denoising Diffusion Implicit Models (DDIMs) (Song et al., 2020b) have made strides in reducing NFEs for diffusion-based generative models, significant computational challenges persist.
In contrast to diffusion models, which parameterize the dynamics of noise-to-data transformation, Conditional Flow Matching (CFMs) directly parameterize vector fields governing this transformation (Liu et al., 2022; Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023). CFMs leverage the change of variable theory in statistics to define conditional probability paths, enabling faster convergence during training compared to diffusion models (Lipman et al., 2022). However, generating high-quality samples with CFMs still requires hundreds of NFEs, making the process computationally expensive (Dao et al., 2023) and limiting the extension of CFM models to time-series data modeling, as NFEs scale with sequence length.
To address these challenges, we introduce a novel method to enhance the efficiency of CFMs, allowing for longer sampling steps without compromising the fidelity of the target distribution. Our approach draws inspiration from Hamiltonian Monte Carlo (HMC) algorithms (Neal, 2012a; Neal et al., 2011), which utilize the conservation properties of Hamiltonian dynamics to couple target and momentum variables. The key contributions of our work are as follows:
@@
    We propose Implicit Dynamical Flow Fusion (IDFF), which integrates a momentum term into the vector field of conditional flow models. Similar to how HMC reduces the number of samples required to reach the target distribution (Neal, 2012a; Neal et al., 2011), IDFF reduces the NFE by a factor of 10 compared to traditional CFMs while maintaining identical marginal distributions. Additionally, we introduce a training objective that directly aligns generated samples with actual data rather than solely focusing on matching conditional flows.
@@
    We validate IDFF through extensive experiments on image generation using datasets such as CIFAR-10 and CelebA. Additional results on ImageNet-64, CelebA-HQ, LSUN-Church, and LSUN-Bedroom are presented in Figure 1. Our models achieve a superior balance between computational cost and sample quality compared to both traditional diffusion models (e.g., DDPM (Ho et al., 2020), DDIM (Song et al., 2020b)) and current CFMs.
@@
    We demonstrate that reducing NFEs makes IDFF practical for other generation tasks, such as time-series modeling. We showcase IDFF’s effectiveness in capturing complex dynamical systems, including 3D attractors, molecular dynamics (MD), and sea surface temperatures (SST). Our results indicate that IDFF surpasses other methods in sample quality for MD and SST datasets while maintaining computational efficiency, with NFE $\leq$ 5 for each generated sample.
Figure 1: Image generation using IDFF (with NFE=10) applied to multiple datasets: CIFAR10, CelebA-64, ImageNet-64, LSUN-Bedroom, LSUN-Church, and CelebA-HQ. The quality of the generated images demonstrates IDFF’s ability to capture and reproduce realistic visuals across varying levels of complexity and resolution. Additional samples and further analysis can be found in Appendix A.5.
@@ 2 Background
Score-based generative models.
[…]
@@
Score-based models (Song and Ermon, 2019) represent a broad class of diffusion models that describe a continuous-time stochastic process characterized by a stochastic process $\mathbf{x}_{t}$, which is governed by the following Itô stochastic differential equation (SDE):
$d\mathbf{x}_{t}=\mathbf{f}(\mathbf{x}_{t},t)dt+g(t)d\mathbf{w},$ (1)
@@
where $t\in[0,1]$, $\mathbf{f}(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, $g(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient of $\mathbf{x}_{t}$, and $\mathbf{w}\in\mathbb{R}^{d}$ is a standard Wiener process.
@@
The time-reversed version of this diffusion process, derived from the Fokker–Planck equations (Song and Ermon, 2019), is also a diffusion process. This reverse-time SDE is defined as:
$d\mathbf{x}_{t}=\left(\mathbf{f}(\mathbf{x}_{t},t)-g(t)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right)dt+g(t)d\bar{\mathbf{w}},$ (2)
@@
where $\bar{\mathbf{w}}$ is a standard Wiener process. Equation 2 produces the same marginal distributions as the forward diffusion process defined by Equation 1.
Equation 2 must be solved to generate samples from this model. The score-matching approach simplifies this process by redefining $\mathbf{f}(\mathbf{x}_{t},t)=h(t)\mathbf{x}_{t}$, where $h:\mathbb{R}\rightarrow\mathbb{R}$ typically assumes an affine form, such as $\mathbf{f}(\mathbf{x}_{t},t)=\mathbf{x}_{t}$. Moreover, $g(\cdot)$, which governs the noise intensity added to the process, often follows a linear or exponential schedule in time.
This reformulation enables a simplification of solving Equation 2 by focusing solely on learning $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t})$ for all $t$. With this approach, and initializing with random noise, e.g., $\mathbf{x}_{0}\sim\mathcal{N}(0,\mathbf{I})$, the process gradually converges to the data distribution $p_{\text{data}}$, reaching it at $t=1$.
(Conditional) Flow Matching (CFM).
Flow-matching (FM) transforms a simple prior distribution into a complex target distribution. FM is defined by a time-dependent vector field $\mathbf{v}_{t}$, which governs the dynamics of an ordinary differential equation (ODE):
$\frac{d}{dt}(\mathbf{x}_{t})=\mathbf{v}_{t}(\mathbf{x}_{t})$ (3)
The solution to this ODE, represented by $\phi(\mathbf{x}_{t})$, evolves along the vector field $\mathbf{v}_{t}$ from time 00 to time $1$, producing $\mathbf{x}_{1}\sim p_{1}$. As samples move along the vector field $\mathbf{v}_{t}$, $p_{t}({\mathbf{x}}_{t})$ changes over time, described by the continuity equation.
$\frac{\partial p_{t}}{\partial t}=-\nabla\cdot(p_{t}\mathbf{v}_{t})$ (4)
[…]
If the probability path $p_{t}({\mathbf{x}}_{t})$ and the vector field ${\mathbf{v}}_{t}({\mathbf{x}}_{t})$ are given, and if $p_{t}({\mathbf{x}}_{t})$ can be efficiently sampled, we can then define a parameterized vector field $\hat{{\mathbf{v}}}_{t}({\mathbf{x}}_{t};\bm{\theta})$ using a neural network with weights $\bm{\theta}$. This parameterized vector field approximates $\mathbf{v}_{t}(\mathbf{x}_{t})$, and our goal is to learn how to generate samples from the ODE defined in equation 3. The parameters of the neural network are trained using the following objective: $\mathcal{L}_{\text{FM}}(\bm{\theta}):=\mathbb{E}_{t\sim\mathcal{U}(0,1),{\mathbf{x}}_{t}\sim p_{t}({\mathbf{x}}_{t})}\|\hat{\mathbf{v}}_{t}({\mathbf{x}}_{t};\bm{\theta})-\mathbf{v}_{t}({\mathbf{x}}_{t})\|^{2}$ to solve the ODE defined in equation 3. Here, $\mathbf{v}_{t}({\mathbf{x}}_{t})$ can be defined as $\mathbf{v}_{t}({\mathbf{x}}_{t}):=\mathbb{E}_{q({\mathbf{z}})}\left[\frac{\mathbf{v}_{t}({\mathbf{x}}_{t}|{\mathbf{z}})p_{t}({\mathbf{x}}_{t}|{\mathbf{z}})}{p_{t}({\mathbf{z}})}\right]$ with some conditional variable ${\mathbf{z}}$ with distribution $q({\mathbf{z}})$; refer to Theorem 3.1 in Tong et al. (2023a) for more details.
However, this learning objective becomes intractable for general source and target distributions. To mitigate this, we can focus on cases where the conditional probability paths $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$ and vector fields ${\mathbf{v}}_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{1})$ associated with $p_{t}(\mathbf{x}_{t})$ and ${\mathbf{v}}_{t}({\mathbf{x}}_{t})$ are known and have simple forms (Tong et al., 2023b). In such cases, we can recover the vector field ${\mathbf{v}}_{t}({\mathbf{x}}_{t})$ using an unbiased stochastic objective known as the Conditional Flow Matching (CFM) loss, defined as:
$\mathcal{L}_{\text{CFM}}(\bm{\theta}):=\mathbb{E}_{t,\mathbf{x}_{1},\mathbf{x}_{t}}\left\|\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}|\mathbf{x}_{1};\bm{\theta})-\mathbf{v}_{t}(\mathbf{x}_{t})\right\|^{2}$ (5)
where $t\sim\mathcal{U}[0,1]$, $\mathbf{x}_{1}\sim p_{\text{data}}$, and $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$. The training objective in Equation 5 ensures that the marginalized vector field $\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}|\mathbf{x}_{1};\bm{\theta})$, denoted as $\hat{\mathbf{v}}_{t}(\mathbf{x}_{t})$, generates $p_{t}(\mathbf{x}_{t})$, similar to FM models.
CFMs vs Diffusion Models.
@@
CFMs share similarities with diffusion models; both rely on defining continuous probability paths over time. However, diffusion models use stochastic processes to transform data distributions, typically governed by an SDE. However, CFMs utilize a vector field to directly map an initial prior distribution to the target distribution via a deterministic process (Equation 4). Diffusion models, such as denoising score matching models, rely on stochastic diffusion paths to approximate the data distribution. CFMs bypass this stochastic process by constructing a conditional vector field and sampling from the ODE defined in Equation 3.
Optimal Transport CFMs (OT-CFMs).
@@
The CFM objective defined in Equation 5 appears simple. However, it is nearly intractable due to the absence of prior knowledge regarding an appropriate mapping that links between the initial ($p_{0}$) and target ($p_{\text{data}}$) distributions. To address this challenge, optimal transport (OT) (Tong et al., 2020) is used to couple the initial and target distributions. This approach seeks a mapping from $p_{0}$ to $p_{1}$ that minimizes the displacement cost between these two distributions, resulting in flows that can be integrated accurately. This CFM variant is known as Optimal Transport CFMs (OT-CFMs), and the conditional vector field corresponds to this is $\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})=\frac{\mathbf{x}_{1}-\mathbf{x}_{t}}{1-t}$. While OT improves training efficiency in CFMs, traditional sampling methods for CFMs require over a hundred function evaluations (NFE) to produce high-quality samples (Tong et al., 2023a). The lack of flexibility in sampling steps during training further limits their ability to efficiently generate samples, often resulting in a high NFE.
Additionally, the OT-CFM vector field, which can be reformulated as $\hat{\mathbf{v}}_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})=\mathbf{x}_{1}-\mathbf{x}_{0}$ Pooladian et al. (2023), focuses on exactly transporting all points ${\mathbf{x}}_{0}$ to ${\mathbf{x}}_{1}$ along straight-line paths. This implies that all points along the path from 00 to $t$ means that the transport map shares the same regression weights when training with the $\mathcal{L}_{\text{CFM}}(\bm{\theta})$ objective. However, this uniform weighting across the trajectory may overlook that points closer to $t=1$ need finer attention to detail and, consequently, a more complex flow; this may limit the model’s ability to effectively represent complex structures in the final generated samples. Furthermore, the deterministic sample generation in CFMs, ${\mathbf{x}}_{t}={\mathbf{x}}_{t-\Delta t}+\Delta t\hat{{\mathbf{v}}}_{t}({\mathbf{x}}_{t};\bm{\theta})$ (see Figure 2.B for visualization), may restrict the flexibility of CFM, as it prevents the introduction of stochasticity that could otherwise enhance the diversity and richness of generated samples. This might manifest in a lack of expressiveness when generating fine-grained or high-dimensional data. These limitations underscore the need for new approaches to sample diversity and NFE of CFMs. Our paper tackles these limitations by proposing an alternative vector field that better balances deterministic transport with elements of stochasticity to improve sample quality.
@@
Figure 2: A) Comparison of trajectory sampling between IDFF and OT-CFMs: The figure displays 4096 final samples generated by IDFF. As shown, IDFF takes larger steps toward the target distribution, guided by the momentum term. While OT-CFM follows a nearly straight path to reach the target distribution, it requires a higher number of function evaluations (NFEs). B) OT-CFMs sampling process. C) IDFF sampling process. In this process, $\hat{\mathbf{x}}_{1}(.)$ approximates the data sample ${\mathbf{x}}_{1}$, and $\hat{\epsilon}(.)$ approximates the scores associated with $\bm{\xi}_{t}$. For further details, see Algorithm 2. The key difference between IDFF and OT-CFMs is that instead of directly generating the vector field with, IDFF generates $\hat{\mathbf{x}}_{1}(.)$ and $\hat{\epsilon}(.)$, and then reconstructs the vector field, following it. Additionally, IDFF uses a Gaussian distribution to sample ${\mathbf{x}}_{t}$ rather than calculating it deterministically.